OpenStudy (anonymous):

If 3x+y=17 and x+2y=5,then 2x-y/4= A)3 B)5 C)7 D)9 E)11

OpenStudy (anonymous):

subtract 3x+y=17 (-) x+2y=5 -------------------- 2x-y =? ......... u proceed

OpenStudy (anonymous):

someone explain more better please